# Understanding In-Context Learning with a Pelican Soup Framework

### v1

@@ 6 Real-world Evidence
We inspect whether LMs can do ICL with pronouns well because pronouns are reference words frequently associated with different meaning and our framework suggests that LMs learn ICL ability by modeling the association between reference words and their meaning. We thus experiment with the template [input]”, [verbalizer] thought. and use “he”, “she” as the verbalizers. We follow the setup in Min et al. (2022a) and compare the accuracy of the binary classification tasks using GPT2-Large.
Table 3 shows that this task-agnostic template with pronouns is competitive with those task-specific templates. This contradicts with the belief that only larger models can do in-context learning with task-irrelevant verbalizer Wei et al. (2023). It suggests that modeling reference-meaning may indeed contribute LMs’ ICL ability.
@@ 7 Related Work
Since Brown et al. (2020) discovered large language models’ in-context learning ability, some theoretical works have attempted to explain how language models acquire this ability. Based on a hidden Markov model (HMM) assumption on the language generation process, Xie et al. (2022) suggested that in-context learning is an implicit Bayesian inference process. Hahn & Goyal (2023) defined the generation process with Compositional Attribute Grammar, which is weaker than the HMM assumption, explaining the in-context learning ability with the minimum description length. Zhang et al. (2023) assumed a more general latent variable model. Arora & Goyal (2023) analyze the emergence of skills based on the scaling law (Hoffmann et al., 2022). While their analysis assumes a set of skills as the atomic elements of NLP tasks, our framework is based on a set of atom concepts.
@@
There were also many empirical studies on the in-context learning ability. Some works focused on the effect of the instruction (Webson & Pavlick, 2022; Lampinen et al., 2022; Jang et al., 2023), while some focused on the examples in the demonstration (Liu et al., 2022; Lu et al., 2022; Sorensen et al., 2022; Min et al., 2022b; Yoo et al., 2022; Ye et al., 2023; Chang & Jia, 2023; Ye et al., 2023; Wang et al., 2023b; Kossen et al., 2023). Shin et al. (2022) found that not all training corpora led to in-context learning ability. Prystawski & Goodman (2023) used synthetic data to suggest that the pretraining dataset’s locality structure contributes to the reasoning steps’ effectiveness. Wang et al. (2023a) studied the reasoning steps in chain-of-thought. Akyürek et al. (2024) formulated ICL as learning a formal language from demonstrations and benchmarked model families.
@@
Some previous work studied in-context learning as a meta-learning-like problem (Chen et al., 2022). Some works focused on the relationships between in-context learning and optimization algorithms (Garg et al., 2022; von Oswald et al., 2022; Akyürek et al., 2023; Fu et al., 2023; Guo et al., 2023). Some works inspected the mechanism of ICL in transformer models (Hendel et al., 2023; Bietti et al., 2023; Todd et al., 2023; Shen et al., 2023; Bai et al., 2023). Chan et al. (2022) studied the properties of dataset distribution that could contribute to the in-context learning ability. Li et al. (2023) provided generalization bounds based on the stability of Transformer models and the distance of downstream tasks. Compared to these works, we focus on how the pretraining data in natural language contributes to the in-context learning ability.
8 Conclusion
@@
In this work, we propose a framework that instantiates a latent variable model for analyzing ICL. Compared with previous works (Xie et al., 2022; Zhang et al., 2023), our latent model better reflects the complexity of language. Under our framework, we establish an ICL loss bond that provides insights into the effect of the choice of verbalizers, downstream task distribution, and instruction tuning. We further provide extensions to the framework that allow us to analyze the generalization to unseen tasks using the results by Arora & Goyal (2023) and relate our bound to the function description length discussed by Hahn & Goyal (2023).
We then present a toy setup, Calcutec, as a concrete instance of our framework. Calcutec mimics the distribution shifts LMs need to overcome to perform ICL. Our empirical results show that transformers trained with autoregressive loss can generalize under these distributional shifts. We additionally use an addition task to show that larger models generalize better, potentially explaining why scaling up model size helps LM performance. Finally, we experiment with real-world NLP tasks and show that GPT-Large can do ICL with pronouns, providing evidence supporting our framework. We expect that our toy datasets can facilitate future research on ICL.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgements
We appreciate Joshua Robinson, Ollie Liu, and Xinyan Velocity Yu for providing valuable feedback on this paper.
@@ Appendix B Proof of Theorem 2
Let $S_{t}=x_{1},r_{2},x_{2},r_{2},\cdots,x_{t},r_{t}$.
$\Pr(z,g|S_{t},K)=$ $\frac{\Pr(S_{t}|z,g,K)\Pr(z|K)}{\sum_{z}\Pr(S_{t}|z,g,K)\Pr(z,g|K)}$
$=$ $\frac{\Pr(z^{\prime},g|K)\prod_{i=1}^{t}\Pr(x_{t},r_{t},d|z,g,S_{t-1},K)}{\sum_{z^{\prime},g}\Pr(z,g|K)\prod{i=1}^{t}\Pr(x_{t},r_{t},d|z^{\prime},g,S_{t-1},K)}$
$P(x_{t+1},r_{t+1},d|S_{t},K)=$ $\sum_{z,g}\Pr(x_{t+1},r_{t+1},d|z,S_{t},K)\Pr(z,g|S_{t},K)$
$=$ $\frac{\sum_{z,g}\Pr(z,g|K)\prod_{i=1}^{t+1}\Pr(x_{t},r_{t},d|z,g,S_{t-1},K)}{\sum_{z^{\prime}}\Pr(z,g|K)\prod_{i=1}^{t}\Pr(x_{t},r_{t},d|z^{\prime},g,S_{t-1},K)}.$
@@
$-\sum_{t=0}^{T}\log\Pr(x_{t+1},r_{t+1},d|S_{t},K)=$ $-\sum_{t=0}^{T}\begin{aligned} \Big{(}&\log\sum_{z,g}\Pr(z,g|K)\prod_{i=1}^{t+1}\Pr(x_{t},r_{t},d|z,g,S_{t-1},K)\\ &\quad-\log\sum_{z,g}\Pr(z,g|K)\prod_{i=1}^{t}\Pr(x_{t},r_{t},d|z,g,S_{t-1},K)\Big{)}\end{aligned}$
$=$ $-\log\sum_{z,g}\Pr(z,g|K)\prod_{i=1}^{T+1}\Pr(x_{t},r_{t},d|z,g,S_{t-1},K)$
$\leq$ $-\log\Pr(z^{*},\dot{g}|K)\prod_{i=1}^{T+1}\Pr(x_{t},r_{t},d|z^{*},\dot{g},S_{t-1},K)$
$=$ $-\log\Pr(z^{*},\dot{g}|K)-\sum_{i=1}^{T+1}\log\Pr(x_{t},r_{t},d|z^{*},\dot{g},S_{t-1},K)$
$=$ $-\log\Pr(z^{*},\dot{g}|K)-\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},\dot{g},S_{t-1},K)-\sum_{i=1}^{T}\log\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K).$
@@
$-\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1},K)$
$\leq$ $-\frac{1}{T}\left(\log\Pr(z^{*},\dot{g}|K)+\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},\dot{g},S_{t-1},K)+\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K)}{\Pr(x_{t}|S_{t-1},K)}\right)$
$+\frac{1}{T}\sum_{i=1}^{T}\log\Pr(d|r_{t},x_{t},S_{t-1},K).$
$\leq$ $-\frac{1}{T}\left(\log\Pr(z^{*},\dot{g}|K)+\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},\dot{g},S_{t-1},K)+\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K)}{\Pr(x_{t}|S_{t-1},K)}\right)$
Appendix C Proof of Corollary 3
@@
$\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K)}{\Pr(x_{t}|S_{t-1},K)}=$ $\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},K)}{\Pr(x_{t}|S_{t-1},K)}$
$=$ $\sum_{x_{1},x_{2},\cdots,x_{T}}\Pr(x_{t}|z^{*},\dot{g},K)\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},K)}{\Pr(x_{t}|S_{t-1},K)}$
$=$ $\mathrm{KLD}(\Pr(x_{t}|z^{*},\dot{g},K)||\Pr(x_{t}|S_{t-1},K))\geq 0$
Appendix D The Connection between $P(r_{t}|x_{t},z^{*},\ddot{g})$ and Function Description Length by Hahn & Goyal (2023)
@@
$\Pr(r_{t}|x_{t},z^{*},\ddot{g})=\sum_{\pi\in\Pi}\Pr(\pi,r_{t}|x_{t},z^{*},\ddot{g})\Pr(\text{$\pi$ is dropped}),$
@@
$\log\Pr(r_{t}|x_{t},z^{*},\ddot{g})=$ $\log\sum_{\pi\in\Pi}\Pr(\pi,r_{t}|x_{t},z^{*},\ddot{g})\Pr(\text{$\pi$ is dropped})$
$\geq$ $\log\Pr(\pi^{*},r_{t}|x_{t},z^{*},\ddot{g})\Pr(\text{$\pi^{*}$ isdropped})$
@@
Then we can discuss the connection between $\Pr(r_{t}|x_{t},z^{*},\ddot{g})$ and the function description length by Hahn & Goyal (2023). We can see the dropped reasoning steps in $\pi^{*}$ as the hidden (tree) structure that maps $x_{t}$ to $r_{t}$ as the derivation tree $\tau_{\phi}$ in the bound of Hahn & Goyal (2023). The length of the reasoning steps thus corresponds to the description length of the derivation tree $\mathrm{D}(\tau_{\phi})$.
A major difference between the bound by Hahn & Goyal (2023) and our bound is that their bound has $\mathrm{D}(\tau_{\phi})$ constant to $T$ while our bound has $\sum_{t}\log\Pr(r_{t}|x_{t},z^{*},\ddot{g})$, which potentially grows proportionally to $T$. The cause of this difference is that, Hahn & Goyal (2023) assumes a structure that repetitively applies a function mapping in a document, and the number of repetition is independent to the complexity of the function mapping. In comparison, our framework does not make this assumption.
Appendix E Detailed Gengeration Process of the LM Training Data in Calcutec

### v2

@@ 8 Real-world Evidence
We inspect how likely that LMs learn ICL by modeling the expression-meaning association as our framework suggests. We specifically inspect a small model GPT2-Large and the setting in which we use pronouns as verbalizers. Because pronouns are reference words frequently associated with different meanings, we expect that even a small model can do ICL well with pronouns. To do so, we experiment with the template [input]”, [verbalizer] thought. and use “he”, “she” as verbalizers. We follow the setup in Min et al. (2022a) and compare the accuracy of binary classification tasks, including SST-2 Socher et al. (2013), CR Hu and Liu (2004), MR Pang and Lee (2005), and Subj Pang and Lee (2004), using GPT2-Large.
Table 4 shows that this task-agnostic template with pronouns is competitive with those task-specific templates. This aligns with our conjecture that the high frequency of pronouns in the training set allows smaller LMs to acquire the ICL capability. It also shows that, unlike what Wei et al. (2023) claims, not only larger models can do in-context learning with task-irrelevant verbalizers. On the other hand, using task-irrelevant verbalizers “foo” and “bar” has lower performance (which aligns with the observation of Wei et al. (2023)). This may be because learning to perform ICL with those low-frequency words requires more training data, and thus a larger model size.
@@ 9 Related Work
Since Brown et al. (2020) discovered large language models’ in-context learning ability, some theoretical works have attempted to explain how language models acquire this ability. Based on a hidden Markov model (HMM) assumption on the language generation process, Xie et al. (2022) suggested that in-context learning is an implicit Bayesian inference process. Hahn and Goyal (2023) defined the generation process with Compositional Attribute Grammar, which is weaker than the HMM assumption, explaining the in-context learning ability with the minimum description length. They also studied the compositionality of natural language tasks with function compositions. Zhang et al. (2023) assumed a more general latent variable model. Arora and Goyal (2023) analyze the emergence of skills based on the scaling law (Hoffmann et al., 2022). While their analysis assumes a set of atomic skills for NLP tasks, our framework is based on a set of atom concepts.
@@
There were also many empirical studies on the in-context learning ability. Some works focused on the effect of the instruction (Webson and Pavlick, 2022; Lampinen et al., 2022; Jang et al., 2023), while some focused on the examples in the demonstration (Liu et al., 2022; Lu et al., 2022; Sorensen et al., 2022; Min et al., 2022b; Yoo et al., 2022; Ye et al., 2023; Chang and Jia, 2023; Ye et al., 2023; Wang et al., 2023b; Kossen et al., 2023). Shin et al. (2022) found that not all training corpora led to in-context learning ability. Prystawski and Goodman (2023) used synthetic data to suggest that the pretraining dataset’s locality structure contributes to the reasoning steps’ effectiveness. Wang et al. (2023a) studied the reasoning steps in chain-of-thought. Akyürek et al. (2024) formulated ICL as learning a formal language from demonstrations and benchmarked model families.
@@
Some previous work studied in-context learning as a meta-learning-like problem (Chen et al., 2022). Some works focused on the relationships between in-context learning and optimization algorithms (Garg et al., 2022; von Oswald et al., 2022; Akyürek et al., 2023; Fu et al., 2023; Guo et al., 2023). Some works inspected the mechanism of ICL in transformer models (Hendel et al., 2023; Bietti et al., 2023; Todd et al., 2023; Shen et al., 2023; Bai et al., 2023). Chan et al. (2022) studied the properties of dataset distribution that could contribute to the in-context learning ability. Li et al. (2023b) provided generalization bounds based on the stability of Transformer models and the distance of downstream tasks. We instead focus on how the pretraining data in natural language contributes to the ICL learning ability.
10 Conclusion and Future Work
@@
In this work, we propose a framework that explains how linguistic phenomena in the training corpus lead to LLMs’ ICL and instruction-following capability. Compared with previous works Xie et al. (2022); Zhang et al. (2023), our latent model better reflects the complexity of language. By introducing the notion of knowledge base and logic system, our framework provides insights into how LLMs can generalize from pretraining to downstream tasks, instantiating a setup compatible with the assumptions made by Arora and Goyal (2023). We also relate our bound to the function description length discussed by Hahn and Goyal (2023).
Our framework illuminates a few possible directions for improving LLMs:
  1. 1.
    Our work highlights the importance of learning the interrelation between meanings. As previous works have shown that the language modeling objective is inefficient for this purpose (Allen-Zhu and Li, 2023; Chiang et al., 2024), we suggest that developing a more sophisticated learning algorithm is crucial.
  2. 2.
    Our theory illustrates how linguistic studies on parsing functions can inform LLM research, offering a new theoretical foundation for analyzing LLM generalization.
  3. 3.
    The experimental results of our arithmetic task show that Transformer models can generalize to unseen prompts by modeling the intermediate step-by-step reasoning process. This may be related to the success of the symbolic chain-of-thought distillation (Li et al., 2023a; Hsieh et al., 2023; Shridhar et al., 2023). Investigating and strengthening the mechanism can improve the efficiency of LM training.
11 Limitations
A limitation of our framework is that, as most theoretical studies do, we simplify the real-world scenario to draw insights. One simplification we make is that, we do not take the noise in LLMs’ training data into account. As we may need to make more assumption on the noise to establish a generic theoretical result, we leave it for future study. Another simplification is that, we assume that the language model can perfectly model the distribution of natural language (i.e., only generates outputs satisfying the assumptions in §3.1, especially Assumption 3.2). However, it is unlikely to be the case in practice. On the one hand, the training data may not cover all the test cases. On the other hand, LLMs may not perfectly generalize from the training set. We need to make more assumptions on the training/test data distribution and/or have a deeper understanding on how deep learning models generalize to alleviate this assumption. Therefore, we deem this out of the scope of this paper.
We also note some limitations of our experiments. As commonly seen in theoretical work, our experiment setup in §7.1 is a simplification of natural language. Our intention is to isolate and illustrate specific theoretical mechanisms without confounding variables introduced by the full complexity of natural language. The experiment in §8 shows the correlation between better ICL performance and the use of pronouns as verbalizer. Although this serves as evidence aligning with our theoretical framework, more studies are required to establish the causal relationship.
@@ Appendix B Proof of Theorem 1
Let $S_{t}=x_{1},r_{2},d,x_{2},r_{2},d\cdots,x_{t},r_{t},d$.
$\Pr(z|S_{t})$
$=$ $\frac{\Pr(S_{t}|z)\Pr(z)}{\sum_{z}\Pr(S_{t}|z)\Pr(z)}$
$=$ $\frac{\Pr(z)\prod_{i=1}^{t}\Pr(x_{i},r_{i},d|z,S_{i-1})}{\sum_{z^{\prime}}\Pr(z^{\prime})\prod_{i=1}^{t}\Pr(x_{i},r_{i},d|z^{\prime},S_{i-1})}$
$P(x_{t+1},r_{t+1},d|S_{t})$
$=$ $\sum_{z}\Pr(x_{t+1},r_{t+1},d|z,S_{t})\Pr(z|S_{t})$
$=$ $\frac{\sum_{z}\Pr(z)\prod_{i=1}^{t+1}\Pr(x_{i},r_{i},d|z,S_{i-1})}{\sum_{z^{\prime}}\Pr(z)\prod_{i=1}^{t}\Pr(x_{i},r_{i},d|z^{\prime},S_{i-1})}.$
@@
$-\sum_{t=0}^{T}\log\Pr(x_{t+1},r_{t+1},d|S_{t})$
$=$ $\begin{aligned} &-\sum_{t=0}^{T}\Big(\\ &\log\sum_{z}\Pr(z)\prod_{i=1}^{t+1}\Pr(x_{i},r_{i},d|z,S_{i-1})\\ &-\log\sum_{z}\Pr(z)\prod_{i=1}^{t}\Pr(x_{i},r_{i},d|z,S_{i-1})\\ &\Big)\end{aligned}$
$=$ $-\log\sum_{z}\Pr(z|K)\prod_{i=1}^{T+1}\Pr(x_{i},r_{t},d|z,S_{i-1})$
$\leq$ $-\log\Pr(z^{*})\prod_{i=1}^{T+1}\Pr(x_{i},r_{t},d|z^{*},S_{i-1})$
$=$ $-\log\Pr(z^{*})$
$-\sum_{i=1}^{T+1}\log\Pr(x_{i},r_{i},d|z^{*},S_{i-1})$
$=$ $-\log\Pr(z^{*})$
$-\sum_{i=1}^{T}\log\Pr(r_{i},d|x_{i},z^{*},S_{i-1})$
$-\sum_{i=1}^{T}\log\Pr(x_{i}|z^{*},S_{i-1}).$
@@
$-\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1})$
$\leq$ $-\frac{1}{T}\bigg(\log\Pr(z^{*})$
$+\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},S_{t-1})$
$+\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},S_{t-1})}{\Pr(x_{t}|S_{t-1})})$
$+\frac{1}{T}\sum_{i=1}^{T}\log\Pr(d|r_{t},x_{t},S_{t-1}\bigg).$
$\leq$ $-\frac{1}{T}\bigg(\log\Pr(z^{*})$
$+\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},S_{t-1})$
$+\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},S_{t-1})}{\Pr(x_{t}|S_{t-1})}\bigg)$
Appendix C Bounding ICL for Generation
We can extend the ICL bound shown in Theorem 1 to generation tasks. When using ICL for generation, in the demonstration part of the prompt, people usually include a separator between the input and the output. For example, if the task is to translate a sentence to English, people may write the demonstration in the format “[input] should be converted to [output].” This “should be converted to” is an expression that can be associated to the “meaning” of the task, namely “should be mapped to a translation in English.” The intuition of the theorem is that, a language model may be able to uncover this association through the demonstration examples via the same mechanism as how it perform ICL for classification tasks.
Theorem C.1 (Average ICL Loss for Generation).
Let $z^{*}$ represent the association between a separator $\xi$ and a task descriptions such that for any input $x$, $P(x,\xi,r|z^{*})>0$ only if $r$ is one of the correct outputs as specified by the task. Let $K$ be the constraints used for decoding, and $\dot{g}$ be the event where a document follows certain formats. Let $R_{t}$ be the set of correct outputs for $x_{t}$, $S_{t}=\{x_{1},\xi,r_{2},d,x_{2},\xi,r_{2},\cdots,x_{t},\xi,r_{t},d|r_{1}\in R_{1},r_{2}\in R_{2},\cdots,r_{t}\in R_{t}\}$. We have for any integer $T>0$, the average cross-entropy loss of ICL is bounded as:
$\begin{split}&-\frac{1}{T}\sum_{t=0}^{T}\log\Pr(R_{t}|x_{t},\xi,S_{t-1})\\ &\leq\begin{aligned} &-\frac{1}{T}\log\Pr(z^{*})\\ &-\frac{1}{T}\sum_{i=1}^{T}\log\Pr(R_{t},d|x_{t},\xi,z^{*},S_{t-1})\\ &-\frac{1}{T}\sum_{i=1}^{T}\log\frac{\Pr(x_{t},\xi|z^{*},S_{t-1})}{\Pr(x_{t},\xi|S_{t-1})}\end{aligned}\end{split}$ (4)
Proof.
If we see $\xi$ as a part of the $x$’s in the proof for Theorem 1, then following the steps in §B, we can have Eq. 4. ∎
Appendix D Proof of Corollary 2
@@
$\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},S_{t-1},K)}{\Pr(x_{t}|S_{t-1},K)}$
$=$ $\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},K)}{\Pr(x_{t}|S_{t-1},K)}$
$=$ $\sum_{x_{1},x_{2},\cdots,x_{T}}\Pr(x_{t}|z^{*},K)\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},K)}{\Pr(x_{t}|S_{t-1},K)}$
$=$ $\mathrm{KLD}(\Pr(x_{t}|z^{*},K)||\Pr(x_{t}|S_{t-1},K))\geq 0$
Appendix E The Connection between $P(r_{t}|x_{t},z^{*})$ and Function Description Length by Hahn and Goyal (2023)
@@
$\begin{split}&\Pr(r_{t}|x_{t},z^{*})\\ &=\sum_{\pi\in\Pi}\Pr(\pi,r_{t}|x_{t},z^{*})\Pr(\text{$\pi$ is dropped}),\end{split}$
@@
$\log\Pr(r_{t}|x_{t},z^{*})$
$=$ $\log\sum_{\pi\in\Pi}\Pr(\pi,r_{t}|x_{t},z^{*})\Pr(\text{$\pi$ is dropped})$
$\geq$ $\log\Pr(\pi^{*},r_{t}|x_{t},z^{*})\Pr(\text{$\pi^{*}$ is dropped})$
@@
Then we can discuss the connection between $\Pr(r_{t}|x_{t},z^{*},\ddot{g})$ and the function description length by Hahn and Goyal (2023). We can see the dropped reasoning steps in $\pi^{*}$ as the hidden (tree) structure that maps $x_{t}$ to $r_{t}$ as the derivation tree $\tau_{\phi}$ in the bound of Hahn and Goyal (2023). The length of the reasoning steps thus corresponds to the description length of the derivation tree $\mathrm{D}(\tau_{\phi})$.
A major difference between the bound by Hahn and Goyal (2023) and our bound is that their bound has $\mathrm{D}(\tau_{\phi})$ constant to $T$ while our bound has $\sum_{t}\log\Pr(r_{t}|x_{t},z^{*},\ddot{g})$, which potentially grows proportionally to $T$. The cause of this difference is that, Hahn and Goyal (2023) assumes a structure that repetitively applies a function mapping in a document, and the number of repetition is independent to the complexity of the function mapping. In comparison, our framework does not make this assumption.
Appendix F Details of the Calcutec Experiment
F.1 Generation Process of the LM Training Data in Calcutec